OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is -11 and the last term is 859,375, the common ratio being -5? a) -143,231 b) -36,047 c) 144,177 d) 716,144

OpenStudy (anonymous):

@praxer Your answer was up for a moment and then disappeared for me =c

OpenStudy (anonymous):

do you have your math book handy? look in the example are of the section for "geometric series" equation

OpenStudy (anonymous):

*example area

OpenStudy (praxer):

follow the second formula... not the first one.. Okay ????? since r < 1 so the second one if the r > 1 than the first one.. here S is the sum and r is the common term and the mu is the first term and n is the no. of terms.....

OpenStudy (anonymous):

@DemolisionWolf I take online courses and couldn't find the equations in my lessons, @praxer thank you very much for the help! Both times you've answered have helped me greatly ^^.

OpenStudy (praxer):

$$\huge \dfrac{-11*(1-(-5)^8)}{1+5}$$ the answer is 716144. Do check the calculations. Once again.......

OpenStudy (anonymous):

Why does the first -5 stay negative and the bottom -5 goes positive?

OpenStudy (praxer):

in the formula in the bottom it is 1-r since r is negative so 1-(-5)=1+5..... Got it ?..... :)

OpenStudy (praxer):

I skipped one step :P

OpenStudy (anonymous):

oh....I think i understand. Thank you!

OpenStudy (praxer):

$$\huge \frak WELCOME....... \ddot\smile $$